A basketball is dropped from the top of a dam. If the ball is dropped from a height of 221 meters, how long does it take to hit the ground?



1
Expert's answer
2020-12-14T12:15:13-0500

We can find the time that the basketball takes to hit the ground from the kinematic equation:


"y=\dfrac{1}{2}gt^2,""t=\sqrt{\dfrac{2y}{g}}=\sqrt{\dfrac{2\cdot 221\ m}{9.8\ \dfrac{m}{s^2}}}=6.7\ s."

Answer:

"t=6.7\ s."